OpenStudy (anonymous):

What is the freezing point of a solution that contains 21.2g of urea,CO(NH2)2, in 265mL water,H2O? Assume a density of water of 1.00g/mol

OpenStudy (anonymous):

you are going to need to calculate the molality of your soln, then figure out the ionization factor, then plug it into the freezing point despersion formula.

OpenStudy (anonymous):

soln???

OpenStudy (anonymous):

solution but when i'm in a hurry i write it soln, another common one i might do is rxn for reaction.

OpenStudy (anonymous):

ohkayy. i REALLY need help with this problem because i have atest tomorrow & i have no idea how to do it -_-

OpenStudy (anonymous):

Ok and i'm willing to help you figure this out but you have to figure out the things I tell you to, like first how many ions will Co(NH2)2 break into?

OpenStudy (anonymous):

okay, thanks. it breaks into 3. Co, N & H

OpenStudy (anonymous):

ok remember that number, now we need to find the molality of the soln. \[molality=\frac{mole of solute}{kg solvent}\]

OpenStudy (anonymous):

ΔT = (Kf)(m)(i), This is equation we are trying to find the different parts for

OpenStudy (anonymous):

its actually 4..C,O,N,H

OpenStudy (anonymous):

it's 3; Co and 2 NH, but see if you can get the molality figured out and I will try to find the constant we need

OpenStudy (anonymous):

its CO not Co

OpenStudy (anonymous):

the molalrity is C=12. O=16 N2= 14 X 2= 28 H4=1X 4=4 so the malalrity is 60

OpenStudy (anonymous):

that looks like the molar mass to me, now i would divide the amount of urea you have by the molar mass to get the moles of solute which will be in your soln.

OpenStudy (anonymous):

so 4/60 right?

OpenStudy (anonymous):

no it would be 21.2g/ 60g =? divide that number by .265 and we will get the m value m=1.33

OpenStudy (anonymous):

\[\Delta T=1.86 * 1.33\]

OpenStudy (anonymous):

\[\Delta T=2.48\] now subtract that from waters normal freezing point and you get -2.48 degree celcius

OpenStudy (anonymous):

yay thanks so much

OpenStudy (anonymous):

Good luck on your test tomorrow, remember to figure out if you solute is ionic or molecular, because this changes up the equation slightly. Practice your molality they are big for these types of problems.